# Fourier Neural Operators for Learning Dynamics in Quantum Spin Systems

### v1

@@ I.2.2 Time-domain Architecture
(a) Refer to caption
(b) FNO n Input Type Fidelity[$\frac{3}{2}\pi,\frac{5}{2}\pi$] Fidelity[$\frac{5}{2}\pi,\frac{23}{2}\pi$] 4 Random 0.$9$$\pm$ 0.00e-03 0.9996 $\pm$ 3.00e-03 4 Low Energy 0.$9$$\pm$ 0.00e-03 0.9992 $\pm$ 1.00e-03 8 Random 0.9996 $\pm$ 2.43e-05 0.9978 $\pm$ 7.81e-05 8 Low Energy 0.9997 $\pm$ 3.83e-05 0.9179 $\pm$ 9.55e-03
(c) U-Net
n Input Type Fidelity[$\frac{3}{2}\pi,\frac{5}{2}\pi$] Fidelity[$\frac{5}{2}\pi,\frac{23}{2}\pi$]
4 Random 0.9966 $\pm$ 6.52e-05 0.9580 $\pm$ 8.00e-03
4 Low Energy 0.$9$$\pm$ 0.00e-03 0.7876 $\pm$ 6.10e-02
8 Random 0.9945 $\pm$ 5.37e-05 0.9790 $\pm$ 8.42e-05
8 Low Energy 0.9991 $\pm$ 3.59e-05 0.1492 $\pm$ 6.50e-03
Figure 3: (a) Prediction of temporal dynamics in an 8-qubit Heisenberg spin system using the time-domain architecture, with FNO and U-Net. The input time interval is denoted as $[0,\frac{3}{2}\pi]$, as illustrated in Figure 1 d), while the extrapolated time refers to future time predictions beyond the trained range. (b) The results present the mean fidelity for both the training time (i.e., the time intervals the FNO encountered during training) and the extrapolated time [$\frac{5}{2}\pi,\frac{23}{2}\pi$] for systems with 4 and 8 qubits, across different input types. (c) The results show the mean fidelity using U-Net for similar inputs and time intervals.
@@
where $m=\frac{3}{2}T/\Delta t$ defines the number of of equidistant time-steps in the interval $[0,\frac{3}{2}T]$. For training, we choose $T=\pi$ and $\Delta t=\pi/10$. For the positional embedding $\operatorname{Embed_{\mathrm{time}}}$, we use the values of two sinusoidal functions applied to the time-grid in order to assist the FNO in learning temporal patterns. The output of the FNO is given as
@@
where $S_{[T,\frac{5}{2}T]}$ denotes the wavefunction evolved over the (discretized) time interval $[T,\frac{5}{2}T]$, where we overlap the time interval $[T,\frac{3}{2}T]$ with the one of the input I in Eq. (9). This setup facilitates smoother and more accurate learning, as the FNO can leverage the temporal continuity and patterns in the evolved wavefunction data. Consequently, the output includes the future time interval $[\frac{3}{2}T,\frac{5}{2}T]$, unseen during training. We use two different sets of previously described inputs, i.e., random wavefunctions and low-energy wavefunctions. While the basis states does not necessarily need to be ordered as in the previous architecture, we maintain the order for consistency.
We also perform extrapolated time prediction by applying the model on its predicted time intervals to predict unseen future time intervals, up to $t=\frac{23}{2}T$, as seen in Figure 3. The fidelity metrics for both the training intervals and the extrapolated times are also reported in Figure 3, demonstrating that even with substantial predictions into the future, the FNO achieves exceptionally high fidelities for systems of both 4 and 8 qubits. Additionally, we compare the performance of the FNO with that of a deep neural network, specifically a U-Net [45]. Our evaluation demonstrates that the FNO outperforms the U-Net, especially in the extrapolated time regime, as seen in Figure 3. Moreover, even at moderate system sizes (8 qubits), the FNO achieves an 6.71x speedup compared to the exact unitary-based method in predicting the dynamics of random inputs up to $t=\frac{23}{2}T$, with only a negligible fidelity reduction of $0.04\%$.
Additionally, we performed evaluations on the output time interval $[T,\frac{5}{2}T]$ using a finer grid with width $\Delta t=\pi/100$ for a 4-qubit system with random inputs. For the FNO, the resulting fidelity of $0.9999$ was identical to that obtained on the coarser training discretization. While we can also apply the U-Net on the finer grid, it is not discretization-agnostic, leading to an error rate of $4.78\%$ in the expected fidelity. To further demonstrate the FNO’s capability to achieve zero-shot super-resolution, we used an input interval of $[0,\frac{3}{2}T]$ with $T=5\pi$ and $\Delta t=\pi/2$ for 4 qubits and evaluated the corresponding output interval on a finer grid with $\Delta t=\pi/20$. The FNO successfully predicted the finely discretized output with an error rate of just $0.04\%$, whereas the U-Net demonstrated a significantly higher error rate of $51.70\%$ compared to the coarse grid fidelity.
@@ I.3 Learning Dynamics using Hamiltonian Observables
Figure 4: Mean Relative Error (MRE) for predicted Hamiltonian observables in an 8-qubit system using FNO, utilizing the time-domain architecture. The model is trained on the input time interval $[0,\frac{3}{2}\pi]$ as shown in Figure 1 d) and predicts the output interval $[\frac{3}{2}\pi,\frac{5}{2}\pi]$. Additionally, it extrapolates future unseen dynamics $[\frac{2}{2}\pi,\frac{7}{2}\pi]$, which is twice the length of the output interval. For extrapolated time predictions on ground truth (GT), the MRE is $6.46\%$. This is particularly significant in quantum simulations, where the primary objective is to extend these simulations. Additionally, based on the time interval the model was trained on, the MRE for extrapolated time predictions is $11.18\%$ denoted by orange dashes. We additionally calculate the mean MSE loss for the training time interval as $3.096\times 10^{-6}\pm 2.497\times 10^{-6}$ and for future time predictions over the same interval on ground truth as $4.212\times 10^{-6}\pm 3.333\times 10^{-6}$.
@@
Hamiltonian observables refer to the individual terms (observables) from the time-evolution generator (Hamiltonian) of the quantum wavefunction. These observables are obtained using the wavefunction by calculating the expectation values for the individual observable terms. For example, in the Ising model, the observables could include terms like $\sigma_{z}\sigma_{z}$ or $\sigma_{z}$ while in the Heisenberg model, they could include $\sigma_{a}\sigma_{a}$ (where $a\in\{x,y,z\}$) and $\sigma_{z}$. Using observables from the Hamiltonian, instead of the full wavefunction, as inputs and outputs for the FNO greatly enhances its scalability, as there are approximately $\text{poly}(n)$ terms in the former and $2^{n}$ terms in the latter. This compression would allow us to train and make predictions with quantum systems that are too large or long to simulate directly, e.g., by using data from large quantum devices or wide tensor networks, from which we could then infer times longer than the devices’ coherence time or past the tensor network’s tractable depth, respectively. However, predicting future dynamics based on these partial observables presents a significant challenge, particularly without explicit knowledge of the underlying Hamiltonian.
@@
To evaluate the FNO’s capabilities, we first considered the Ising Hamiltonian as described in Eq. (2), where $J_{z}$ and $h$ are randomly generated. Importantly, in this study, we focus on using random wavefunction states to calculate the expectation values of these observables, making the learning and prediction tasks for the FNO maximally general and, as a result, challenging.
@@
For the observables, we exclusively utilize the time-domain architecture. In this setup, the input consists of predefined observables and their time-evolved intervals. We provide a detailed description of the specific observables used in Section III. This input structure mirrors the previously described input I in Eq. (9), but instead of wavefunction basis states $S$, it includes observable quantities. This setup presents a substantial challenge for the FNO, as it must learn to predict future dynamics based on only partial information about the quantum state.
@@
We further extrapolate to the future time interval $[\frac{5}{2}T,\frac{7}{2}T]$, effectively doubling the dynamics captured within the output time interval $[\frac{3}{2}T,\frac{5}{2}T]$. This feature is particularly advantageous for quantum computing, where the limited coherence time of error-corrected qubits imposes constraints on the length of computations. Additionally, it can also be integrated to take observable data from tensor network methods and extrapolate it to timescales that would be computationally difficult for it to calculate due to high tensor train rank. Given that these observables present challenges for the FNO to learn effectively, we extrapolate future times based on the ground truth of the predicted time interval. This approach helps minimize unnecessary errors, as the ground truth is typically known. As seen in Figure 4, with a system of 8 qubits we predict future observables with a relative error of $6.46\%$, even when extending the prediction to double the train time horizon on ground truth.
@@ II Discussions
We have presented two FNO architectures that are capable of not only learning the time-evolutions of quantum systems, but also of extrapolating these evolutions to later times. The energy-domain architecture is more compact, requiring fewer computational inputs, and hones in on key dynamics by prioritizing slow quantum transitions. The time-domain architecture is agnostic to the discretization of the time interval, making it highly versatile for obtaining the output state of quantum evolution at arbitrary times. Both methods demonstrate the ability of FNOs to not only carry out quantum state evolution, but to learn the underlying time-evolution operator itself, which constitutes a key accomplishment in the ubiquitous and challenging task of solving the Schrödinger equation.
Of foremost interest is the application of the time-domain FNO using only Hamiltonian observables as input and output. This difficult learning task requires that the FNO learn and conduct future inference on partial information, using only $\sim\text{poly}(n)$ expectation values rather than the full $2^{n}$-component wavefunction. Under this architecture, the FNO could feasibly be trained using measurements from noisy quantum devices or data from shallow tensor networks, then infer to later time-scales that are otherwise out of reach of contemporary techniques as they would require either more coherent quantum computers or intractably large tensor networks.
@@ III Methods
We provide details on the wavefunction ordering protocol used in Section I.2.1, the types of Pauli strings utilized in Section I.3, and the specifics of the training data, and FNO configurations for Figures 2, 3, and 4.
@@
In Section I.3, for a system of 8 qubits, we use a set of 48 observables that includes all nearest-neighbor interactions $XX$, $YY$, and $ZZ$ interactions across all qubit pairs. Additionally, the set includes single-qubit interactions $X$, $Y$ and $Z$. We focus exclusively on quantities exceeding a certain threshold (e.g., greater than $10^{-2}$) to avoid including less significant values that might inflate the relative error. Given that these observables are the expectation values of Pauli strings, we employ Mean Squared Error (MSE) and Mean Relative Error (MRE) as our loss metrics to evaluate the model’s performance.
@@
In Figure 4, we use 18,000 training data points with 48 Pauli strings for the 8-qubit system, utilizing 4 layers and retaining 7 modes.
Acknowledgements.
F.S. acknowledges support from the Caltech Summer Undergraduate Fellowship. J.B. acknowledges support from the Wally Baer and Jeri Weiss Postdoctoral Fellowship. A.A.’s work is supported in part by the Bren endowed chair, the ONR (MURI grant N00014-18-12624), and the AI2050 Senior Fellow Program at Schmidt Sciences.

### v2

@@ I.2.2 Time-domain Architecture
Figure 3: (a) Prediction of temporal dynamics in a four-qubit Heisenberg spin system using the time-domain architecture in Section I.2.2. The model is trained on the time interval $0,3\pi/2]$ as illustrated in Figure 1d, while the extrapolated time refers to future predictions beyond the trained range. Performance for both the random and low-energy wavefunction initial states using the FNO is benchmarked against a U-Net baseline. The shaded regions around each curve indicate one standard deviation. The FNO accurately captures the underlying operator and extrapolates reliably to later times, while the U-Net exhibits poor extrapolation, especially for low-energy input states. (b) Corresponding results for an eight-qubit Heisenberg spin system.
@@
where $m=\frac{3}{2}T/\Delta t$ defines the number of of equidistant time-steps in the interval $0,\frac{3}{2}T]$. For training, we choose $T=\pi$ and $\Delta t=\pi/10$. For the positional embedding $\operatorname{Embed}_{\mathrm{time}}$, we use a sinusoidal encoding to facilitate the FNO in learning temporal patterns. Given a discretized time grid with $m$ equidistant points, indexed by $l\in\{0,\dots,m-1\}$, the positional embedding is constructed using a single cosine–sine pair. With two embedding channels and a fixed frequency $\omega=1$, each time index $l$ is mapped as
$\operatorname{Embed}_{\mathrm{time}}(l)=\big[\cos(\omega l),\;\sin(\omega l)\big].$ (10)
This results in a positional embedding tensor of shape $(1,2,m)$, where the two channels correspond to the cosine and sine components evaluated at each time step. By mapping time onto a smooth periodic representation, the embedding allows the model to capture the relative ordering of temporal points while preserving continuity across the discretized interval [24, 8]. Although the approach can be generalized to multiple frequencies by increasing the number of channels, in this work we use the minimal two-channel formulation. The output of the FNO is given as
@@
where $S_{[T,\frac{5}{2}T]}$ denotes the wavefunction evolved over the (discretized) time interval $T,\frac{5}{2}T]$, where we overlap the time interval $T,\frac{3}{2}T]$ with the one of the input I in Eq. (9). This setup facilitates smoother and more accurate learning, as the FNO can leverage the temporal continuity and patterns in the evolved wavefunction data. Consequently, the output includes the future time interval $\frac{3}{2}T,\frac{5}{2}T]$, unseen during training. We use two different sets of previously described inputs, i.e., random wavefunctions and low-energy wavefunctions. While the basis states does not necessarily need to be ordered as in the previous architecture, we maintain the order for consistency.
We also perform extrapolated time prediction by applying the model on its predicted time intervals to predict unseen future time intervals, up to $t=\frac{23}{2}T$, as seen in Figure 3. The fidelity metrics for both the training intervals and the extrapolated times are also reported in Figure 3, demonstrating that even with substantial predictions into the future, the FNO achieves exceptionally high fidelities for systems of both 4 and 8 qubits. Additionally, we compare the performance of the FNO with that of a deep neural network, specifically a U-Net [46]. Our evaluation demonstrates that the FNO outperforms the U-Net, especially in the extrapolated time regime, as seen in Figure 3. Moreover, even at moderate system sizes (8 qubits), the FNO achieves an 6.71x speedup compared to the exact unitary-based method in predicting the dynamics of random inputs up to $t=\frac{23}{2}T$, with only a negligible fidelity reduction of $0.04\%$. This speedup does not take into account the training cost, which for a moderate size of 8 qubits is approximately $2$s per epoch over 200 training epochs.
Additionally, we performed evaluations on the output time interval $T,\frac{5}{2}T]$ using a finer grid with width $\Delta t=\pi/100$ for a 4-qubit system with random inputs. For the FNO, the resulting fidelity of $0.9999$ was identical to that obtained on the coarser training discretization. While we can also apply the U-Net on the finer grid, it is not discretization-agnostic, leading to an error rate of $4.78\%$ in the expected fidelity. To further demonstrate the FNO’s capability to achieve zero-shot super-resolution, we used an input interval of $0,\frac{3}{2}T]$ with $T=5\pi$ and $\Delta t=\pi/2$ for 4 qubits and evaluated the corresponding output interval on a finer grid with $\Delta t=\pi/20$. The FNO successfully predicted the finely discretized output with an error rate of just $0.04\%$, whereas the U-Net demonstrated a significantly higher error rate of $51.70\%$ compared to the coarse grid fidelity.
@@ I.3 Learning Dynamics using Hamiltonian Observables
(a) Refer to caption
n $\mathrm{MRE}(24\pi/10)$ $\mathrm{MRE-GT}(5\pi/2)$ $\mathrm{MRE-GT}(31\pi/10)$ $\mathrm{MRE-GT}(34\pi/10)$ 8 0.0675 $\pm$ 0.0305 0.0242 $\pm$ 0.0114 0.0815 $\pm$ 0.0484 0.1057 $\pm$ 0.0573 20 0.0460 $\pm$ 0.0208 0.0459 $\pm$ 0.0176 0.0515 $\pm$ 0.0150 0.0704 $\pm$ 0.0247
@@
Figure 4: (a) Mean Relative Error (MRE) for predicted Hamiltonian observables in a 20-qubit system using FNO, utilizing the time-domain architecture. The model is trained on the input time interval $0,\frac{3}{2}\pi]$ as shown in Figure 1 d) and predicts the output interval $\frac{3}{2}\pi,\frac{5}{2}\pi]$. Additionally, it extrapolates future unseen dynamics $\frac{5}{2}\pi,\frac{7}{2}\pi]$, which is twice the length of the output interval. The MRE is $5.8\%$ over the extrapolated time predictions on ground truth (GT). This is particularly significant in quantum simulations, where the primary objective is to extend these simulations. Additionally, based on the time interval the model was trained on, the MRE for extrapolated time predictions is $9.6\%$ denoted by orange dashes. We additionally calculate the mean MSE loss for the training time interval as $2.16\times 10^{-9}\pm 2.15\times 10^{-9}$ and for future time predictions over the same interval on ground truth as $5.34\times 10^{-9}\pm 4.13\times 10^{-9}$. (b) The results present the MRE at specific time steps for both the train time (i.e., the time intervals that FNO encountered during training) and the extrapolated time $\frac{5}{2}\pi,\frac{7}{2}\pi$] for systems with 8 and 20 qubits for hamiltonian observables.
@@
Hamiltonian observables refer to the individual operator terms that appear in the time-evolution generator (the Hamiltonian) of the system. In practice, we do not use these operators themselves as inputs to the model. Rather, we compute their expectation values with respect to the evolving quantum state and use those quantities as the actual inputs. For example, in the Ising model, the relevant operators include terms such as $\sigma_{z}\sigma_{z}$ and $\sigma_{z}$, while in the Heisenberg model they include $\sigma_{a}\sigma_{a}$ for $a\in\{x,y,z\}$ and $\sigma_{z}$. We evaluate the corresponding expectation values $\langle O_{j}\rangle_{t}=\langle\psi(t)|O_{j}|\psi(t)\rangle$ at each time step, which then form the real-valued input and output channels for the FNO. Using expectation values of Hamiltonian, instead of the full wavefunction, as inputs and outputs for the FNO greatly enhances its scalability, as there are approximately $\text{poly}(n)$ terms in the former and $2^{n}$ terms in the latter. This compression would allow us to train and make predictions with quantum systems that are too large or long to simulate directly, e.g., by using data from large quantum devices or wide tensor networks, from which we could then infer times longer than the devices’ coherence time or past the tensor network’s tractable depth, respectively. However, predicting future dynamics based on these partial observables presents a significant challenge, particularly without explicit knowledge of the underlying Hamiltonian.
@@
To evaluate the FNO’s capabilities, we first considered the Ising Hamiltonian as described in Eq. (2), where $J_{z}$ and $h$ are randomly generated. This introduces variability in the Hamiltonian parameters, ensuring that the learning task is general with respect to the Hamiltonian and hence challenging. Importantly, in this study, we focus on using random wavefunction states to calculate the expectation values of these observables.
@@
For the observable-based inputs, we exclusively utilize only the time-domain architecture. In this configuration, the model receives a fixed set of observable expectation values over the chosen time interval. A detailed specification of the observables appears in Section III. This input structure mirrors the previously described input I in Eq. (9), but instead of $2^{n}$ wavefunction basis states $S$, it includes a substantially smaller set of observable expectation values. We use the same input time interval, $[0,3T/2]$, with $T=\pi$ and time step $\Delta t=\pi/10$, as described in Section I.2.2. This setup presents a substantial challenge for the FNO, as it must learn to predict future dynamics based on only partial information about the quantum state.
@@
For our experiments, we consider system sizes of 8 and 20 qubits. For the larger system (20 qubits), the time-evolved expectation-value data is generated using the CUDA-Q library with Trotterization, since exact simulation at these scales is computationally intractable. We further extrapolate to the future time interval $\frac{5}{2}T,\frac{7}{2}T]$, effectively doubling the dynamics captured within the output time interval $\frac{3}{2}T,\frac{5}{2}T]$. This feature is particularly advantageous for quantum computing, where the limited coherence time of error-corrected qubits imposes constraints on the length of computations. Additionally, it can also be integrated to take observable data from tensor network methods and extrapolate it to timescales that would be computationally difficult for it to calculate due to high tensor train rank. Given that these observables present challenges for the FNO to learn effectively, we extrapolate future times based on the ground truth of the predicted time interval. This approach helps minimize unnecessary errors, as the ground truth is typically known. As seen in Figure 4, the FNO maintains consistent and accurate performance even as the system size increases substantially. For 8- and 20-qubit systems, we predict future observables with relative errors of $6.46\%$ and $5.8\%$, respectively, even when extending the prediction to twice the training time horizon on ground truth data. The model exhibits no degradation in accuracy with increasing system size, and its extrapolation performance remains stable across all settings. This strongly demonstrates that FNOs can scale effectively to larger quantum many-body systems while preserving error rates in both the training regime and long-time extrapolation. It is also noteworthy that the required volume of training data does not scale in proportion to the exponential growth of the Hilbert space. For instance, we use 18,000 training samples for the 8-qubit system and only 22,500 for the 20-qubit system. This modest increase, despite the dramatic growth in system dimensionality, underscores the data efficiency of the FNO architecture and further illustrates its practical scalability for large-scale quantum dynamical prediction.
Once trained, FNO enables extremely fast inference of quantum dynamics. For a representative example involving 1500 time steps, FNO inference requires 7.1 s in total, whereas GPU-accelerated Trotterized time evolution implemented in CUDA-Q requires approximately 85 s per time step on a Grace Hopper system and 180 s per time step on an NVIDIA A100 GPU. This corresponds to an inference-time speedup of approximately $1.8\times 10^{4}$ (Grace Hopper) to $3.8\times 10^{4}$ (A100) compared to numerical time-evolution methods. Although this does not include the training cost of the FNO, this cost is incurred only once. For the largest system considered (20 qubits), training requires approximately 40 s per epoch over 1700 epochs, corresponding to a total training time of about 19 hours using an A100 GPU. This cost is substantially reduced for smaller systems, such as 8 qubits, where the training time per epoch is approximately 22 s. While training the FNO entails a substantial one-time computational cost that grows with system size, this cost can be amortized across multiple subsequent inference tasks, such as computing dynamics for new initial states. In contrast, numerical time-evolution methods must be rerun independently for each such trajectory, resulting in a computational cost that grows linearly with the number of simulations.
@@ II Discussions
We have presented two FNO architectures that are capable of not only learning the time-evolutions of quantum systems, but also of extrapolating these evolutions to later times. The energy-domain architecture is more compact, requiring fewer computational inputs, and hones in on key dynamics by prioritizing slow quantum transitions. The time-domain architecture is agnostic to the discretization of the time interval, making it highly versatile for obtaining the output state of quantum evolution at arbitrary times. However, this approach does not scale efficiently, since the Hilbert space dimension of the wavefunction increases exponentially with the number of qubits. Nevertheless, it remains useful for faster evaluation as, for an 8-qubit system, it yields a 6.716.71x speedup in evaluation compared to exact solvers. Furthermore, because the FNO learns the underlying operator and considers the full wavefunction as input, it can reliably extrapolate to approximately 1010x the training time, a regime that is computationally demanding for exact solvers and inaccessible to tensor-network methods due due to increasing entanglement. Both methods demonstrate the ability of FNOs to not only carry out quantum state evolution, but to learn the underlying time-evolution operator itself, which constitutes a key accomplishment in the ubiquitous and challenging task of solving the Schrödinger equation.
Of foremost interest is the application of the time-domain FNO using only Hamiltonian observables as input and output. This difficult learning task requires that the FNO learn and conduct future inference on partial information, using only $\sim\text{poly}(n)$ expectation values rather than the full $2^{n}$-component wavefunction. Notably, our results demonstrate that the FNO maintains consistent performance as system size increases substantially. As shown in Figure 4, for 8- and 20-qubit systems, the FNO predicts future observables with relative errors of 6.46% and 5.8%, respectively, even when extrapolating to twice the training time horizon on ground truth data. Furthermore, it requires only a modest increase in the number of training samples to scale efficiently to larger quantum many-body systems.
These results establish FNO as a promising computational framework for exploring novel quantum phenomena and simulating complex quantum many-body systems. Its favorable scalability, capacity for long-time extrapolation, and data-efficient learning enable access to regimes that are challenging for conventional numerical approaches. In particular, the FNO could be explored as a data-driven framework to learn the dynamics of a broad class of systems, including interacting spin chains, fermionic lattice models, and disordered quantum systems that are otherwise computationally intractable. Within this architecture, the FNO may be trained on measurement data obtained from noisy intermediate-scale quantum devices or on data generated by shallow tensor-network simulations, and subsequently leveraged to predict system dynamics at later time scales. Such predictions would otherwise necessitate either highly coherent quantum hardware or tensor-network representations of prohibitive bond dimension.
@@ III Methods
We provide details on the low-energy states generation in Section I.2, wavefunction ordering protocol used in Section I.2.1, the types of Pauli strings utilized in Section I.3, and the specifics of the training data, and FNO configurations for Figures 2, 3, and 4.
We generate the low-energy states referenced in Section I.2 by computing the energies of all $2^{n}$ computational-basis states under the system Hamiltonian and select the $M$ configurations with the lowest energies. A low-energy wavefunction is constructed by assigning complex amplitudes only to these $M$ modes, sampled independently from a uniform distribution, while all remaining basis-state amplitudes are set to zero. The resulting state is normalized to have unit $L^{2}$ norm. In our numerical settings, we use $M=4$ for the 4-qubit system and $M=50$ for the 8-qubit system.
@@
In Section I.3, for a system of 8 qubits, we use a set of 48 observables that includes all nearest-neighbor interactions $XX$, YYYY, and $ZZ$ interactions across all qubit pairs. Additionally, the set includes single-qubit interactions $X$, $Y$ and $Z$. We focus exclusively on quantities exceeding a certain threshold (e.g., greater than $10^{-2}$) to avoid including less significant values that might inflate the relative error. Given that these observables are the expectation values of Pauli strings, we employ Mean Squared Error (MSE) and Mean Relative Error (MRE) as our loss metrics to evaluate the model’s performance.
@@
In Figure 4, we use 18,000 training data points with 48 Pauli strings for the 8-qubit system, utilizing 4 layers and retaining 7 modes. For 20 qubits, we use 22500 training data points with 120 Pauli strings, 4 layers, and retaining 7 modes.
IV acknowledgments
F.S. acknowledges support from the Caltech Summer Undergraduate Fellowship. J.B. acknowledges support from the Wally Baer and Jeri Weiss Postdoctoral Fellowship. A.A.’s work is supported in part by the Bren endowed chair, the ONR (MURI grant N00014-18-12624), and the AI2050 Senior Fellow Program at Schmidt Sciences. We acknowledge the Amazon Web Services (AWS) Advanced Compute team, with particular thanks to Tyler Y. Takeshita and Sebastian Hassinger, for their compute support.
Code availability
The code to reproduce the results reported in this study is available from the corresponding author upon request.